OpenStudy (anonymous):

What is an equation in point slope form for the line perpendicular to y=4x-5 that contains (-10,3)?

OpenStudy (***[isuru]***):

if 2 lines r perpendicular to each other and if the slope of one line is "m" then the slop of the other lone is "(-1/m)" so thw slope of the line u want will be "(-1/4)" then the basic format will be y = mx + b = (-1/4)x + b then consider the coordinates (-10 , 3) and substitute them in the above equations 3 = (-1/4)10 + b b = 3 + 10/4 = 22/4 = 11/2 therefor the line u want is y= (-1/4)x + 11/2 got it :) ?

OpenStudy (anonymous):

Yes, thank you so much. Can you help me with one more?

OpenStudy (***[isuru]***):

if it is within my limit ... i will

OpenStudy (anonymous):

Write the equation for the vertical line that contains point E(6,4).

OpenStudy (***[isuru]***):

a vertical line means that the x value of this is always a constant so the equation is x = 6

OpenStudy (***[isuru]***):



OpenStudy (anonymous):

Thank you.